OpenStudy (anonymous):

What is the simplified form of the quantity 1 over x minus 1 over y all over 1 over x plus 1 over y ? y plus x, over y minus x y minus x, over y plus x x minus y, over x plus y x plus y, over x minus y @mathstudent55

OpenStudy (mathstudent55):

I'll use the equation editor to make sure I get the problem right.

OpenStudy (mathstudent55):

Is this it? \( \huge\dfrac{\frac{1}{x} - \frac{1}{y}}{\frac{1}{x} + \frac{1}{y}} \)

OpenStudy (anonymous):

Yes!

OpenStudy (mathstudent55):

If so, here's an easy way to simplify this fraction. In the numerator there are two denominators, x and y. In the denominator, there are also two denominators, x and y. What is the LCD of x and y?

OpenStudy (anonymous):

My guess is x-y/x+y

OpenStudy (mathstudent55):

I don't guess. I figure it out.

OpenStudy (anonymous):

Lol, I'm just telling you of my guess from when I looked at the problem.

OpenStudy (mathstudent55):

You may very well be correct. Let's find out if your hunch is correct.

OpenStudy (mathstudent55):

Ok, what is the LCD of x and y?

OpenStudy (anonymous):

Alright! Would it be 1?

OpenStudy (mathstudent55):

If you had just 2 fractions like these that you needed to add, you'd need a common denominator. \(\dfrac{2}{x} + \dfrac{3}{y} \) What would you use as a common denominator? Since x and y have no factors in common, you multiply them together to get xy. The least common denominator (LCD) of x and y is xy.

OpenStudy (mathstudent55):

Now back to your problem. You have denominators of x and y. The LCD of the denominators is xy. Let's multiply the numerator and denominator of the main fraction by xy.

OpenStudy (mathstudent55):

\( \huge =\dfrac{xy( \frac{1}{x} - \frac{1}{y} ) }{xy(\frac{1}{x} + \frac{1}{y})} \) Now we distribute the xy in the numeratoir and the denominator: \( \huge = \dfrac{xy( \frac{1}{x} ) - xy(\frac{1}{y} ) }{xy(\frac{1}{x} )+ xy(\frac{1}{y})} \) Now we cancel out the x's and y's that are both in the numerator and denominator: \( \huge = \dfrac{\cancel{x}y( \frac{1}{\cancel{x}} ) - x\cancel{y}(\frac{1}{\cancel{y}} ) }{\cancel{x}y(\frac{1}{\cancel{x}} )+ x\cancel{y}(\frac{1}{\cancel{y}})} \) \(\huge =\dfrac{y - x}{y + x} \)

OpenStudy (mathstudent55):

Now compare the answer above with your guess. The conclusion is be careful with guesses.

OpenStudy (mathstudent55):

Ok, gtg. Bye.

OpenStudy (anonymous):

Yes, now I can see it. Thank you! @mathstudent55

OpenStudy (mathstudent55):

You're welcome.